(S points) A o.5 kg block slides along a horizontal frictionless surface at 2.0 m/s. It is brought to rest by compressing a ve 16. sed long spring of spring constant 800 N/m. How far does the spring get compressed (in cm)? 17. (5 points) A block is released from rest at point P. h2 50 m high, and slides along the frictionless track shown. What is its speed at pointō匹=40 m high? h1 h2 ground level
Answer #1

16. Applying energy conservation,

m v^2 / 2 = k x^2 / 2

0.5 (2^2) = (800)x^2

x = 0.05 m

17. Applying energy conservation,

PEi + KEi = PEf + KEF

m g (50) + 0 = m g (40) + m v^2 /2

v = sqrt(2 x 9.8 x (50 - 40))

v= 14 m/s